OpenStudy (anonymous):

A pair of dice is rolled. What is the probability that the sum is 9 or that the first number is a 2?

OpenStudy (anonymous):

You can get 9 in 4 ways 6+3, 3+6, 5+4 and 4+5 out of the 36 outcomes. So that is 4/36=1/9

OpenStudy (anonymous):

P(A or B) = P(A) + P(B) -P(A and B)

OpenStudy (anonymous):

first number is 2 is 1/6. (one good result from 6)

OpenStudy (anonymous):

P(A or B) = P(A) + P(B) -P(A and B) is right Melissa what is P(A), P(B) and P(A and B) here?